OpenStudy (anonymous):

solve by graphing. x + 2y+4 and 4y - 2x = 24

OpenStudy (crashonce):

graph the intercepts then extend to make the lines

OpenStudy (anonymous):

get the equations into y=mx+b form to help

OpenStudy (danjs):

x + 2y = 4 -2x + 4y = 24 multiply the first equation by 2 ---------------- 2x + 4y = 8 -2x + 4y = 24 add them together, (the x terms cancel out) -------------- 0x + 8y = 32 8y=32 y=4 If y=4, put that into one of the equations, and solve for x=??

OpenStudy (danjs):

The point (x,4) will be the solution, and the intersection of the two lines.

OpenStudy (danjs):

To graph, let x=0, solve for y to get point (0,y) let y=0, solve for x to get point (x,0) Those are 2 points on the line, connect them, do that for both lines.